### v2

# Training-Free Mitigation of Adversarial Attacks on Deep Learning-Based MRI Reconstruction

Abstract
Deep learning (DL) methods, especially those based on physics-driven DL, have become the state-of-the-art for reconstructing sub-sampled magnetic resonance imaging (MRI) data. However, studies have shown that these methods are susceptible to small adversarial input perturbations, or attacks, resulting in major distortions in the output images. Various strategies have been proposed to reduce the effects of these attacks, but they require retraining and may lower reconstruction quality for non-perturbed/clean inputs. In this work, we propose a novel approach for mitigating adversarial attacks on MRI reconstruction models without any retraining. Our framework is based on the idea of cyclic measurement consistency. The output of the model is mapped to another set of MRI measurements for a different sub-sampling pattern, and this synthesized data is reconstructed with the same model. Intuitively, without an attack, the second reconstruction is expected to be consistent with the first, while with an attack, disruptions are present. A novel objective function is devised based on this idea, which is minimized within a small ball around the attack input for mitigation. Experimental results show that our method substantially reduces the impact of adversarial perturbations across different datasets, attack types/strengths and PD-DL networks, and qualitatively and quantitatively outperforms conventional mitigation methods that involve retraining. Finally, we extend our mitigation method to two important practical scenarios: a blind setup, where the attack strength or algorithm is not known to the end user; and an adaptive attack setup, where the attacker has full knowledge of the defense strategy. Our approach remains effective in both cases.
1 Introduction
Magnetic resonance imaging (MRI) is an essential imaging modality in medical sciences, providing high-resolution images without ionizing radiation, and offering diverse soft-tissue contrast. However, its inherently long acquisition times may lead to patient discomfort and increased likelihood of motion artifacts, which degrade image quality. Accelerated MRI techniques obtain a reduced number of measurements below Nyquist rate and reconstruct the image by incorporating supplementary information. Parallel imaging, which is the most clinically used approach, leverages the inherent redundancies in the data from receiver coils  [21, 48, 40], while compressed sensing (CS) utilizes the compressibility of images through linear sparsifying transforms to achieve a regularized reconstruction [41, 3, 23, 30]. Recently, deep learning (DL) methods have emerged as the state-of-the-art for accelerated MRI, offering superior reconstruction quality compared to traditional techniques [24, 35, 4, 53]. In particular, physics-driven DL (PD-DL) reconstruction has become popular due to their improved generalizability and performance [24, 27, 2].
While PD-DL methods significantly outperform traditional MRI reconstruction techniques, these approaches have been shown to be vulnerable to small adversarial perturbations [19, 44], invisible to human observers, resulting in significant variations in the network’s outputs  [6, 24, 64]. Various strategies to improve the robustness of PD-DL networks have been proposed to counter adversarial attacks in MRI reconstruction [14, 10, 29, 50, 37].
However, all these methods require retraining of the network, incurring a high computational cost, while also having a tendency to lead to additional artifacts for clean/non-attack inputs [58].
In this work, we propose a novel mitigation strategy for adversarial attacks on DL-based MRI reconstruction, which does not require any retraining. Our approach utilizes the idea of cyclic measurement consistency [73, 33, 55, 70, 72] with synthesized undersampling patterns. The overarching idea for cyclic measurement consistency is to simulate new measurements from inference results with a new forward model that is from a similar distribution as the original forward model, which should be consistent with the original inference. This idea has been used to improve parallel imaging [73], then rediscovered in the context of DL reconstruction training [33, 55, 72] and uncertainty guidance [70]. In our work, we use this idea in a completely novel direction to characterize and mitigate adversarial attacks. Succinctly, without an attack, reconstructions on synthesized measurements should be cycle-consistent, while with a small adversarial perturbation, there should be large discrepancies between reconstructions from actual versus synthesized measurements. We use this consistency to devise an objective function over the network input to effectively mitigate adversarial perturbations. Our contributions are as follows:
  • We propose a novel mitigation strategy for adversarial attacks, which optimizes cyclic measurement consistency over the input within a small ball without requiring any retraining.
  • We show that the mitigation strategy can be applied in a manner that is blind to the size of the perturbation or the algorithm that was used to generate the attack.
  • Our method readily combines with existing robust training strategies to further improve reconstruction quality of DL-based MRI reconstruction under adversarial attacks.
  • Our results demonstrate effectiveness across various datasets, PD-DL networks, attack types and strengths, and undersampling patterns, outperforming existing methods qualitatively and quantitatively, without affecting the performance on non-perturbed images.
  • Finally, we show that the physics-driven nature of our method makes it robust even to adaptive attacks, where the attacker is aware of the defense strategy and finds the worst-case perturbation that maximize its effectiveness in bypassing the defense algorithm.
2 Background and Related Work
2.1 PD-DL Reconstruction for Accelerated MRI
In MRI, raw measurements are collected in the frequency domain, known as the k-space, using multiple receiver coils, where each coil is sensitive to different parts of the field-of-view. Accelerated MRI techniques acquire sub-sampled data, where the forward model is given as
$\mathbf{y}_{\Omega}=\mathbf{E}_{\Omega}\mathbf{x}+\mathbf{n},$ (1)
where ${\bf y}_{\Omega}\in\mathbb{C}^{M}$ is the measured data across all coils, $\mathbf{E}_{\Omega}\in\mathbb{C}^{M\times N}$ is the forward multi-coil encoding operator, with ${M}>{N}$ in the multi-coil setup [47], $\Omega$ is the undersampling pattern with acceleration rate $R$, $\mathbf{n}$ is measurement noise, and $\mathbf{x}$ is the image to be reconstructed [48]. The inverse problem for this acquisition model is formulated as
$\arg\min_{\bf x}\|{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf x}\|_{2}^{2}+{\cal R}({\bf x})$ (2)
where the first quadratic term enforces data fidelity (DF) with the measurements, while the second term is a regularizer, $\mathcal{R}(\cdot)$. The objective in Eq. (2) is conventionally solved using iterative algorithms [18] that alternate between DF and a model-based regularization term [18].
On the other hand, PD-DL commonly employs a technique called algorithm unrolling [43], which unfolds such an iterative reconstruction algorithm for a fixed number of steps. Here, the DF is implemented using conventional methods with a learnable parameter, while the proximal operator for the regularizer is implemented implicitly by a neural network [24, 2, 53, 25]. The unrolled network is trained end-to-end in a supervised manner using fully-sampled reference data [24, 2] using a loss of the form:
$\arg\min_{\bm{\theta}}\mathbb{E}\Big{[}{\cal L}\big{(}f({\bf z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}}),{\bf x}_{\textrm{ref}}\big{)}\Big{]},$ (3)
where ${\bf z}_{\Omega}={\bf E}_{\Omega}^{H}{\bf y}_{\Omega}$ is the zero-filled image that is input to the PD-DL network [25]; $f(\cdot,\cdot;{\bm{\theta}})$ is the output of the PD-DL network, parameterized by ${\bm{\theta}}$, in image domain; ${\cal L}(\cdot,\cdot)$ is a loss function; ${\bf x}_{\textrm{ref}}$ is the reference image. Unsupervised training that only use undersampled data [61, 62, 4] can be used, though this typically does not outperform supervised learning. In this work, we unroll the variable splitting with quadratic penalty algorithm [18], as in MoDL [2].
Figure 1: Overview of the proposed mitigation strategy. a) If there is an adversarial attack, the k-space corresponding to the reconstructions of MRI data synthesized from previous DL model outputs will be disrupted. b) This idea is used to devise a novel loss function to find a “corrective” perturbation around the input that ensures cyclic measurement consistency.
2.2 Adversarial Attacks in PD-DL MRI Reconstruction
Adversarial attacks create serious challenges for PD-DL MRI reconstruction, where small, visually imperceptible changes to input data can lead to large errors in the reconstructed image [71, 6, 10]. The main idea here is to find the worst-case degradation ${\bf r}$ within a small $\ell_{p}$ ball that will lead to the largest perturbation in the output of the network [6]:
$\arg\max_{{\bf r}:||{\bf r}||_{p}\leq\epsilon}{\cal L}\big{(}f({\bf z}_{\Omega}+{\bf r},{\bf E}_{\Omega};{\bm{\theta}}),f({\bf z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}})\big{)}.$ (4)
We note that this attack calculation is unsupervised, which is the relevant scenario for MRI reconstruction [6, 29, 71], as the attacker cannot know the fully-sampled reference for a given undersampled dataset. In MRI reconstruction, $\ell_{\infty}$ perturbations are commonly used in image domain [71, 6, 37, 29], while $\ell_{2}$ perturbations are used in k-space [50] due to scaling differences between low and high-frequency in Fourier domain. In this work, we concenrate on the well-studied class of $\ell_{\infty}$ adversaries, while examples for the $\ell_{2}$ perturbations are provided in SuppMat Sec. 8.7.
We also note that image domain attacks can be converted to k-space as: ${\bf w}=({\bf E}_{\Omega}^{H})^{\dagger}{\bf r}={\bf E}_{\Omega}({\bf E}_{\Omega}^{H}{\bf E}_{\Omega})^{-1}{\bf r}$, since $M>N$ for multi-coil MRI acquisitions [47]. Note ${\bf w}$ is only non-zero at $\Omega$, and its zerofilled image is ${\bf E}_{\Omega}^{H}{\bf w}={\bf r}$, as expected. In other words, $\ell_{\infty}$ attacks have k-space representations, where only the acquired locations $\Omega$ are perturbed, aligning with the underlying physics of the problem.
Adversarial attacks are typically calculated using a gradient-based strategy [19, 42], where the input is perturbed in the direction of maximal change within the $\ell_{\infty}$ ball. In this study, we use the iterative projected gradient descent (PGD) method [42], as it leads to more drastic perturbations than the single-step fast gradient sign method (FGSM) [19]. Further results with FGSM are included in SuppMat Sec. 8.6. Finally, we note that neural network based attacks have also been used [50], but these are mainly preferred for reduced computation time in training, and often fail to match the degradation caused by iterative optimization-based techniques [28].
2.3 Defense Against Adversarial Attacks in MRI Reconstruction
Incorporation of an adversarial term in the training objective is a common method for robust training, and has been proposed both in the image domain  [29] or k-space [50]. The two common approaches either enforce perturbed outputs to the reference [29]:
$\min_{\bm{\theta}}\>\mathbb{E}\bigg{[}\max_{\|\mathbf{r}\|_{\infty}\leq\epsilon}\mathcal{L}[f_{\bm{\theta}}(\mathbf{z}_{\Omega}+\mathbf{r},{\bf E}_{\Omega};{\bm{\theta}}),\mathbf{x}_{\textrm{ref}})]\bigg{]}$ (5)
or aim to balance normal and perturbed training [50]:
$\min_{\bm{\theta}}\>\mathbb{E}\bigg{[}\max_{\|\mathbf{r}\|_{\infty}\leq\epsilon}$ $\>\mathcal{L}[f_{\bm{\theta}}(\mathbf{z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}}),\mathbf{x}_{\textrm{ref}})]$
$+\lambda\mathcal{L}[f_{\bm{\theta}}(\mathbf{z}_{\Omega}+\mathbf{r},{\bf E}_{\Omega};{\bm{\theta}}),\mathbf{x}_{\textrm{ref}})]\bigg{]},$ (6)
where $\lambda$ is a hyperparameter controlling the trade-off. While such training strategies improve robustness against adversarial attacks, it often comes at the cost of reduced performance on non-perturbed inputs [58]. Another recent method for robust PD-DL reconstruction proposes the idea of smooth unrolling (SMUG) [37]. SMUG [37] modifies denoised smoothing [52], introduces robustness to a regularizer part of the unrolled network. Each unrolled unit of SMUG performs:
${\bf x}_{s}^{i+1}=\arg\min_{\bm{x}}$ $\|{\bf E}_{\Omega}{\bf x}_{s}^{i}-{\bf y}_{\Omega}\|_{2}^{2}$
$+\lambda\|{\bf x}-{\mathbb{E}}_{\eta}[{\cal D}_{\bm{\theta}}({\bm{x}}_{s}^{i}+{\bm{\eta}})]\|_{2}^{2}$ (7)
where ${\cal D}_{\bm{\theta}}$ represents the denoiser network with parameters ${\bm{\theta}}$, and ${\bm{\eta}}\sim{\cal N}({\bf 0},\sigma^{2}{\bf I})$ is random Gaussian noise. During the training, SMUG [37] aims to incoperate $N$ number of Monte Carlo sampling to smooth the denoiser outputs, averaging them before entering the next data fidelity block.
2.4 Why Are Adversarial Attacks Important in DL-Based MRI Reconstruction?
MRI reconstruction pipelines are closed proprietary systems [59], thus it is unlikely that an adversary may successfully inject adversarial perturbations during this process. Nonetheless, adversarial attacks provide a controlled means to understand the worst-case stability and overall robustness of DL-based reconstruction systems [6, 20, 71, 26, 5]. It has been argued both empirically [6] and theoretically [20] that worst-case perturbations are not rare events. In particular, if one samples a new input from a small ball around the worst-case perturbation this still leads to a failed reconstruction [6]. In [20], it is further shown that sampling from Gaussian noise, i.e. the thermal noise model in MRI, leads to such an instability with non-zero probability. Apart from Gaussian noise, there are several other causes of perturbations in an MRI scan, including body motion [69] or hardware issues [31], which are hard to model mathematically, but whose combined effect may lead to similar instabilities for DL-based reconstruction [6]. Similarly, adversarial perturbations, and mitigation algorithms like ours, are critical to understand the robustness of DL reconstruction models in important scenarios, such as performance for rare pathologies [45]. However, these physiological changes are much harder to model and simulate, unlike adversarial attacks, which provide insights into worst-case stability. Finally, we note that our mitigation algorithm is also applicable to unrolled networks in general, and may have applications in broader computational imaging scenarios.
Figure 2: Representative reconstruction results for Cor-PD knee, and Ax-FLAIR brain MRI Datasets at $R=4$. The attack inputs lead to severe disruption in the baseline MoDL reconstruction. Adversarial training improves these, albeit suffering from blurriness. SMUG fails to eliminate the attack. The proposed strategy reduces the artifacts and maintains sharpness. Furthermore it can be combined with the other strategies for further gains (last two columns).
3 Proposed Method for Training-Free Mitigating Adversarial Attacks in PD-DL MRI
3.1 Attack Propagation in Simulated k-space
The idea behind our mitigation strategy stems from cyclic measurement consistency with synthesized undersampling patterns, which has been previously used to improve calibration/training of MRI reconstruction models [73, 33, 55, 70, 72]. For reconstruction purposes, a well-trained model should generalize to undersampling patterns with similar distributions as the acquisition one [35]. To this end, let $\{\Delta_{n}\}$ be undersampling patterns drawn from a similar distribution as $\Omega$, including same acceleration rate, similar underlying distribution, e.g. variable density random, and same number of central lines. Further let
${\bf\tilde{x}}_{\Omega}=f({\bf z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}})$ (8)
be the reconstruction of the acquired data. We simulate new measurements $\tilde{\bf y}_{\Delta_{i}}$ from ${\bf\tilde{x}}$ using the encoding operator ${\bf E}_{\Delta_{n}}$ with the same coil sensitivity profiles as ${\bf E}_{\Omega}$, and let ${\bf z}_{\Delta_{i}}={\bf E}_{\Delta_{i}}^{H}\tilde{\bf y}_{\Delta_{i}}$ be the corresponding zerofilled image. Then the subsequent reconstruction
${\bf\tilde{x}}_{\Delta_{i}}=f({\bf z}_{\Delta_{i}},{\bf E}_{\Delta_{i}};{\bm{\theta}})$ (9)
should be similar to ${\bf\tilde{x}}_{\Omega}$. In particular, we evaluate the similarity over the acquired k-space locations, $\Omega$, as we will discuss in Section 3.2. However, if there is an attack on the acquired lines, either generated directly in k-space or in image domain as discussed in Section 2.2, then this consistency with synthesized measurements are no longer expected to hold, as illustrated in Fig. 1a.
This can be understood in terms of what the PD-DL network does during reconstruction as it alternates between DF and regularization. The DF operation will ensure that the network is consistent with the input measurements, ${\bf y}_{\Omega}$, or equivalently the zerofilled image, ${\bf z}_{\Omega}$. If there is no adversarial attack, we expect the output of a well-trained PD-DL network to be consistent with these measurements, while also showing no sudden changes in k-space [35]. On the other hand, if there is an attack, the output will still be consistent with the measurements, as the attack is designed to be a small perturbation on ${\bf y}_{\Omega}$ or ${\bf z}_{\Omega}$, and thus the small changes on these lines will be imperceptible. Instead, the attack will affect all the other k-space locations $\Omega^{C}$, the complement of the acquired index set, leading to major changes in these lines for the output of the PD-DL network, as depicted in Fig. 1a. Thus, when we resample a new set of indices $\Delta_{i}$ that includes lines from $\Omega^{C}$, under attack the next level reconstruction ${\bf\tilde{x}}_{\Delta_{i}}$ will no longer be consistent with the original k-space data ${\bf y}_{\Omega}$, as measured through $||{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf\tilde{x}}_{\Delta_{i}}||_{2}$. The distortion in the k-space will further propagate as we synthesize more levels of data and reconstruct these, if there is an adversarial attack.
This description of the attack propagation suggests a methodology for detecting such attacks; however, this is not the focus of this paper. As discussed in Section 3.2, the mitigation algorithm can be applied on all inputs, regardless of whether they have been attacked, as the algorithm does not degrade the reconstruction quality if the input is unperturbed. We note that this does not create a major computational burden, since the mitigation algorithm does not change the input in this case, i.e. converges in a single iteration, as shown in SuppMat Sec. 8.4. Thus, to keep the exposition clearer, we focus on mitigation for the reminder of the paper. Nonetheless, a threshold-based detection scheme based on these ideas is presented in SuppMat Sec. 7.
3.2 Attack Mitigation with Cyclic Consistency
Based on the characterization of the attack propagation, we next introduce our proposed training-free mitigation strategy. We note that adversarial attacks of Section 2.2 all aim to create a small perturbation within a ball around the original input. Here the size of the ball specifies the attack strength, the particular algorithm specifies how the attack is generated/propagated within the given ball, and the attack domain/norm specifies the type of $\ell_{p}$ ball and whether it is in k-space or image domain. Succinctly, our mitigation approach aims to reverse the attack generation process, by searching within a small ball around the perturbed input to find a clear input. The objective function for finding this clear input is based on the aforementioned idea of cyclic measurement consistency, and is given as
$\arg\min_{{\bf r^{\prime}}:||{\bf r^{\prime}}||_{p}\leq\epsilon}\mathbb{E}_{\Delta}\bigg{[}\Big{\|}({\bf E}_{\Omega}^{H})^{\dagger}({\bf z}_{\Omega}+{\bf r^{\prime}})-$
${\bf E}_{\Omega}f\Big{(}{\bf E}_{\Delta}^{H}\big{(}{\bf E}_{\Delta}f({\bf z}_{\Omega}+{\bf r^{\prime}},{\bf E}_{\Omega};{\bm{\theta}})+{\bf\tilde{n}}\big{)},{\bf E}_{\Delta};{\bm{\theta}}\Big{)}\Big{\|}_{2}\bigg{]}.$ (10)
Here ${\bf r^{\prime}}$ is a small “corrective” perturbation and ${\bf z}_{\Omega}+{\bf r^{\prime}}$ corresponds to the mitigated/corrected input. Hence the first term, $({\bf E}_{\Omega}^{H})^{\dagger}({\bf z}_{\Omega}+{\bf r^{\prime}})$ corresponds to the minimum $\ell_{2}$ k-space solution that maps to this zerofilled image [71]. The second term is the corresponding k-space values at the acquired indices $\Omega$ after two stages of cyclic reconstruction. Note a small noise term, ${\bf\tilde{n}}$, is added to the synthesized data to maintain similar signal-to-noise-ratio [72, 34]. The expectation is taken over undersampling patterns $\Delta$ with a similar distribution to the original pattern $\Omega$.
The objective function is solved using a reverse PGD approach, which is detailed in Algorithm 1. Note the algorithm performs the expectation in Eq. (3.2) over $K$ sampling pattern $\{\Delta_{k}\}_{k=1}^{K}.$ Notably, our reverse PGD performs a gradient descent instead of the ascent in PGD [42], and includes a projection on to the $\epsilon$ ball to ensure the solution remains within the desired neighborhood.
Algorithm 1 Attack Mitigation
1:$\epsilon,\alpha,{\bf z}_{\Omega}^{pert},{\bf E}_{\Omega},\{{\bf E}_{\Delta_{k}}\}_{k=1}^{K},f(\cdot,\cdot;{\bm{\theta}})$triangleright Inputs
2:Clean version of ${\bf z}_{\Omega}^{pert}$$\triangleright$ Mitigate attack on input
3:${\bf\tilde{z}}_{\Omega}={\bf z}_{\Omega}^{pert}$
4:repeat
5:     Loss $=0$
6:     for $k=1$ to $K$ do
7:         ${\bf\tilde{y}}_{\Omega}=({\bf E}_{\Omega}^{H})^{\dagger}{\bf\tilde{z}_{\Omega}}$
8:         ${\bf\tilde{\tilde{y}}}_{\Omega}=$
9:            ${\bf E}_{\Omega}f\big{(}{\bf E}_{\Delta_{k}}^{H}({\bf E}_{\Delta_{k}}f({\bf\tilde{z}}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}})+{\bf\tilde{n}}),{\bf E}_{\Delta_{k}};{\bm{\theta}}\big{)}$
[…]
11:         Loss = Loss + lossk
12:     end for
13:     ${\bf grad}=\frac{1}{K}\nabla_{{\bf\tilde{z}}_{\Omega}}Loss$
14:     ${\bf\tilde{z}}_{\Omega}={\bf\tilde{z}}_{\Omega}-\alpha\cdot\textrm{sgn}({\bf grad})$
15:     ${\bf\tilde{z}}_{\Omega}=\textrm{clip}_{{\bf z}_{\Omega}^{pert},\epsilon}({\bf\tilde{z}}_{\Omega})$triangleright Projection to $\epsilon$ ball
16:until Converge
Finally, note that this algorithm uses the strength of the attack. However, from a practical viewpoint, it may be beneficial to mitigate the attack without this information, which will not always be available to the end user. In other words, while Algorithm 1 optimizes Sec. 3.2, in the blind case, we additionally optimize its input parameters $\epsilon$ and $\alpha$ jointly. To this end, we propose a blind estimation procedure, where we estimate $\epsilon$ and $\alpha$ iteratively. First, we decrease $\epsilon$ with a linear scheduler for a fixed $\alpha$, starting from a large ball until convergence. Subsequently, we fix $\epsilon$ and decrease $\alpha$ similarly. The alternating process can be repeated, though in practice, one stage is sufficient. Finally, for blind mitigation, we always use $\ell_{\infty}$ ball, even for $\ell_{2}$ attacks in k-space discussed in SuppMat Sec. 8.8, as it contains the $\ell_{2}$ ball of the same radius.
Figure 3: Performance across different attack strengths. Both Adversarial Training and SMUG fail to perform well against attack strengths they were not trained on. In contrast, the proposed training-free mitigation shows good performance across perturbation levels.
Figure 4: Proposed mitigation approach is readily applicable to various PD-DL networks for MRI reconstruction.
3.3 Mitigation Performance on Adaptive Attacks
Recent works have suggested that a good performance on iterative optimization-based attacks may not be a good indicator of robustness, and that the class of adaptive attacks can bypass the defense strategy once the attacker is aware of the defense itself [11]. Adaptive attacks devise a perturbation that not only deceive the baseline (reconstruction) network, but also to bypass the defense strategy [13, 66]. Consequently, adaptive attacks have become the standard when evaluating defenses[57]. Finally, the generation of adaptive attacks also require careful design, as gradient obfuscation phenomenon has been reported in defenses against iterative optimization-based attacks [66, 49, 22, 46, 60, 8], giving a false perception of the model’s security [7].
To generate adaptive attacks, our mitigation algorithm in Algorithm 1 needs to be incorporated into the attack generation objective Eq. 4. To simplify the notation, we define our mitigation function based on Sec. 3.2 as
$g({\bf r^{\prime}};{\bf z}_{\Omega})\triangleq\arg\min_{{\bf r^{\prime}}:||{\bf r^{\prime}}||_{p}\leq\epsilon}\mathbb{E}_{\Delta}\bigg{[}\Big{\|}({\bf E}_{\Omega}^{H})^{\dagger}({\bf z}_{\Omega}+{\bf r^{\prime}})-$
${\bf E}_{\Omega}f\Big{(}{\bf E}_{\Delta}^{H}\big{(}{\bf E}_{\Delta}f({\bf z}_{\Omega}+{\bf r^{\prime}},{\bf E}_{\Omega};{\bm{\theta}})+{\bf\tilde{n}}\big{)},{\bf E}_{\Delta};{\bm{\theta}}\Big{)}\Big{\|}_{2}\bigg{]},$ (11)
which leads to the adaptive attack generation objective:
$\arg\max_{{\bf r}:||{\bf r}||_{p}\leq\epsilon}$ ${\cal L}\big{(}f({\bf z}_{\Omega}+{\bf r},{\bf E}_{\Omega};{\bm{\theta}}),f({\bf z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}})\big{)}$
$+\lambda\>g({\bf r^{\prime};{\bf z}_{\Omega}+{\bf r}}),$ (12)
where the first term ensures finding a perturbation that fools the baseline reconstruction, as in Eq. 4, while the second term integrates our mitigation algorithm. Notably, maximizing the whole objective should lead to a perturbation ${\bf r}$ that not only misleads the baseline reconstruction, but also maximizes the mitigation loss, resulting in an adaptive attack. Further implementation details, including tuning of $\lambda$, and implementing $g(\cdot;\cdot)$ to avoid gradient obfuscation [7] are discussed Sec. 4.2.
4 Experiments
4.1 Dataset Details
Our experiments were performed on publicly available fully-sampled multi-coil knee and brain MRI from fastMRI database [36], which have 15 and 20 receiver coils, respectively. Coronal proton density (Cor-PD) and axial FLAIR (Ax-FLAIR) were used for knee and brain data, respectively. Retrospective equispaced undersampling was applied at acceleration $R=4$ to the fully-sampled data with 24 central auto-calibrated signal (ACS) lines.
4.2 Implementation Details
Baseline Network. The PD-DL network used in this study was a modified version of MoDL [2], unrolled for $10$ steps, where a ResNet regularizer was used [27, 63, 16]. Further details about the architecture and training are provided in SuppMat Sec. 6. All the comparison methods were implemented using this MoDL network to ensure a fair comparison, except for the results on the applicability of our method to different PD-DL networks.
Attack Generation Details. PGD [42] was used to generate the attacks in an unsupervised manner, as detailed earlier for a realistic setup. Additional results with supervised attacks and FGSM are provided in SuppMat Sec. 8.5 and Sec. 8.6, respectively, and lead to the same conclusions. Complex images were employed to generate the attack and gradients, and MSE loss was used.
Comparison Methods. We compared our mitigation approach with existing robust training methods, including adversarial training [29, 50] and Smooth Unrolling (SMUG) [37]. Adversarial training was implemented using Eq. (5[29], while results using Eq. (6[50] is provided in SuppMat Sec. 8.3. Further implementation details for all methods are provided in SuppMat Sec. 6.
Dataset Metric SMUG Adversarial Training (AT) Proposed Method + MoDL / SMUG / AT
Cor-PD PSNR $28.22$ $33.99$ 35.14 / 34.85 / 36.57
SSIM $0.79$ $0.92$ 0.92 / 0.92 / 0.94
Ax-FLAIR PSNR $29.67$ $34.03$ 36.41 / 34.67 / 35.63
SSIM $0.84$ $0.91$ 0.95 / 0.92 / 0.94
Table 1: Population metrics for SSIM/PSNR on all test slices
Cyclic Consistency Details. The synthesized masks $\{\Delta_{k}\}$ were generated by shifting the equispaced undersampling patterns by one line while preserving the ACS lines [72]. In this setting, the number of synthesized masks is $R-1$. For blind mitigation, the linear scheduler for $\epsilon$ started from 0.04, and decreased by 0.01. For this estimated, $\tilde{\epsilon}$, the linear scheduler for $\alpha$ started from $\tilde{\epsilon}$ value and ended at $\tilde{\epsilon}/3.5$. The objective in (3.2) was implemented using normalized $\ell_{2}$ loss where $||({\bf E}_{\Omega}^{H})^{\dagger}({\bf z}_{\Omega}+{\bf r^{\prime}})||_{2}$ was used for normalization.
Adaptive Attack Details. Direct optimization of (3.3) requires the solution of a long computation graph and multiple nested iterations of neural networks. However, this may induce gradient obfuscation [7]. Thus, we followed the exact gradient computation strategy of [13], by unrolling $g(\cdot;\cdot)$ in (3.3) first [65], and then backpropagating through the whole objective. To this end, we define $g_{T}(\cdot;\cdot)$ as the $T$-step unrolled version of $g(\cdot;\cdot)$, and report performance for different values of $T$. We note that larger $T$ leads to more memory requirements, which was handled by checkpointing [13, 32]. Furthermore, the presence of ${\bf\tilde{n}}$ in (3.3) may suggest stochasticity in the system [32]. However, ${\bf\tilde{n}}$ is precalculated for a given input in our mitigation algorithm, and held constant throughout the mitigation. To make the adaptive attack as strong as possible, we pass this information about ${\bf\tilde{n}}$ to the adaptive attack as well, thus letting it have oracle knowledge about it. Finally, for maximal performance of the attack, we first tuned $\lambda$ in (3.3) empirically, then generated the adaptive attacks for $T\in\{10,25,50,100\}$. Details on tuning of $\lambda$ and verification of gradient obfuscation avoidance in our adaptive attacks are given in SuppMat Sec. 9.1 and Sec. 9.2.
4.3 Attack Mitigation Results
This section summarizes all results for attack mitigation, and is sub-divided for each experiment, characterizing our mitigation strategy from different view points.
Performance Across Datasets. We first investigate our approach and the comparison methods on the knee and brain MRI datasets. Fig. 2 shows representative results for $R=4$ for all methods. Baseline PD-DL, MoDL shows a high degree of artifacts under attack. SMUG is able to improve these but still suffers from substantial artifacts. Adversarial training resolves this artifacts, albeit with blurring. The proposed approach successfully mitigates the attacks without any retraining, while maintaining sharpness. We note our method can also be combined with SMUG and adversarial training to further improve their performance. Tab. 1 summarizes the quantitative metrics for all test slices in the datasets, which are consistent with the visual observations.
Performance Across Attack Strengths. We next test the performance of the methods across different attack strengths, $\epsilon\in\{0.01,0.02\}$. Fig. 3 shows the results for both attack strengths using the robust training methods trained with $\epsilon=0.01$ and the proposed mitigation approach. Consistent with Fig. 2, SMUG has artifacts at $\epsilon=0.01$, which gets worse at $\epsilon=0.02$. Similarly, adversarial training struggles at $\epsilon=0.02$, since it was trained at $\epsilon=0.01$, leading to visible artifacts (arrows). On the other hand, our training-free mitigation is successful at both $\epsilon$. This is expected, since no matter how big the $\epsilon$ ball is, the mitigation algorithm explores the corresponding vicinity of the perturbed sample to optimize Sec. 3.2. Further quantitative results are provided in SuppMat Sec. 8.1.
Performance Across Different PD-DL Networks. Next, we hypothesize that our method is agnostic to the PD-DL architecture. To test this hypothesis, we perform our mitigation approach for different unrolled networks, including XPDNet [51], Recurrent Inference Machine [39], E2E-VarNet [54], and Recurrent-VarNet [68]. The implementation details are discussed in the SuppMat. Fig. 4 depicts representative images for clear and perturbed inputs, and our proposed cyclic mitigation results. Overall, all networks show artifacts for perturbed inputs, while our proposed cyclic mitigation algorithm works well on all of them to reduce these artifacts. Further quantitative metrics for these networks are provided in SuppMat Sec. 8.2.
Blind Mitigation Results. These experiments show that in addition to not needing any retraining for mitigation, our approach does not require precise information about how the attack is generated. Fig. 5 shows how the reconstruction improves as we use linear schedulers to find the optimum $(\epsilon,\alpha)$ values. Top row shows the tuning of $\epsilon$ while we keep the step size $\alpha$ constant. After the cyclic loss in Sec. 3.2 stops decreasing, we fix this $\tilde{\epsilon}$ for the projection ball. The bottom row shows the effect of decreasing $\alpha$ for this $\tilde{\epsilon}$ value, from right to left. Further results, including $\ell_{2}$ k-space attacks and quantitative metrics are provided in SuppMat Sec. 8.7 and Sec. 8.8.
Figure 5: Blind mitigation process of finding the optimum $(\epsilon,\alpha)$ parameters and corresponding results. Top row shows $\epsilon$ optimization for a fixed $\alpha$, while the bottom row shows $\alpha$ optimization for the optimum $\epsilon$. This joint optimization leads to a 1.15dB gain over the initial estimate.
Performance Against Adaptive Attacks. Tab. 2 shows the performance of our mitigation algorithm for adaptive attacks with $T\in\{10,25,50,100\}$ unrolls. We note that due to the high computational cost of generating the adaptive attacks with 100 unrolls, we ran the adaptive attack generation and mitigation on a subset of 75 Cor-PD slices, which is why the non-adaptive attack results have lower PSNR than the full test set in Tab. 1. For mitigation of the adaptive attacks, we ran both the unrolled version (used to generate the adaptive attack) and the iterative version (ran until convergence) of Algorithm 1. The average number of iterations for the latter are reported in paranthesis in the last column. Further visual examples for adaptive attack mitigation are provided in SuppMat Sec. 9.
We observe the following: 1) Baseline reconstructions have higher PSNR under adaptive attacks than non-adaptive attacks, as adaptive attacks balance two terms, reducing its focus on purely destroying the reconstruction. This effect increases as $T$ increases in the second term, as expected. 2) For few number of unrolls, adaptive attack degrades performance if mitigated with the unrolled version. For $T<50$, we observe that the unrolled mitigation struggles ($\sim$ 5dB degradation for $T=10$) with the adaptive attack designed for matched number of unrolls. 3) Our mitigation readily resolves adaptive attacks if run until convergence. For large $T\geq 50$ values, the unrolled mitigation also largely resolves the adaptive attacks. 4) Even though adaptive attacks with large $T$ lead to a weaker baseline attack, they degrade the performance of our mitigation more, even though the overall degradation is slight even at $T=100$ (.68dB).
Attack Type #Unrolls (T) Baseline Reconstruction Unrolled Algorithm 1 Iterative Algorithm 1
     Non-adaptive N/A $16.16$ N/A $34.69$
     Adaptive 10 $19.23$ $29.47$ $34.34\>(119\>\text{iters})$
Adaptive 25 $19.32$ $32.79$ $34.16\>(111\>\text{iters})$
Adaptive 50 $19.96$ $33.39$ $34.14\>(105\>\text{iters})$
Adaptive 100 $21.02$ $33.78$ $34.01\>(100\>\text{iters})$
Table 2: PSNR for adaptive attacks on 75 slices from the Cor-PD dataset. Parantheses in the last column indicate the mean iteration for convergence of the iterative algorithm.
These observations all align with the physics-driven design of the mitigation: The PD-DL reconstruction network ends with a data fidelity unit, i.e. the network output is consistent with (perturbed) ${\bf y}_{\Omega}$. Since the attack is a tiny perturbation on data indexed at $\Omega$, it will cause misestimation of lines in $\Omega^{C}$ instead (as in Fig.1a). Our method synthesizes new measurements at $\Delta$ from the latter, and uses it to perform a second reconstruction, which are mapped to $\Omega$ and checked with consistency with ${\bf y}_{\Omega}$. So the only way the mitigation can be fooled is if this cyclic consistency is satisfied, which in turn would indicate that the intermediate recon on $\Omega^{C}$ is good, which effectively mitigates the attack.
4.4 Ablation Study
We perform an ablation study on how many levels of reconstructions are needed for mitigation. In this case, multiple steps of reconstructions and data synthesis can be used to update the loss function in Sec. 3.2. Results, given in SuppMat Sec. 10, demonstrate that enforcing cyclic consistency with multiple levels degrades performance and requires more computational resources. Hence, using 2-cyclic reconstruction stages is the best choice from both performance and computational perspectives.
5 Conclusions
In this study, we proposed a method to mitigate small imperceptible adversarial input perturbations on DL-based MRI reconstructions, without requiring any retraining. We showed our method is robust across different datasets, attack strengths, unrolled networks. Furthermore, our method can be combined with existing robust training methods to further enhance their performance. Additionally, the proposed method can be performed in a blind manner without attack-specific information, such as attack strength or type for further practical applicability. Finally, owing to its physics-based design, our method is robust to adaptive attacks, which have emerged as the standard for robustness evaluation in recent years.
Supplementary Material
In these supplementary materials, we provide implementation details of different methods used in this study (Sec. 6), a complementary attack detection strategy based on the ideas in Sec. 3.1 (Sec. 7), more examples and quantitative results for attack variations discussed in the main text (Sec. 8), details on adaptive attack generation and visual examples (Sec. 9), and details on the ablation study (Sec. 10).
6 Implementation Details
6.1 PD-DL Network Details
MoDL implementation is based on [2], unrolling variable splitting with quadratic penalty algorithm [18] for 10 steps. The proximal operator for the regularizer is implemented with a ResNet [61, 17], and data fidelity term is implemented using conjugate gradient, itself unrolled for 10 iterations [2]. The ResNet comprises input and output convolutional layers, along with 15 residual blocks. Each residual block has a skip connection and two convolutional blocks with a rectified linear unit in between. At the end of each residual block, there is a constant scaling layer [56], and the weights are shared among different blocks [2].
XPDNet implementation is based on [67] and follows [51], which unrolls the primal dual hybrid gradient (PDHG) algorithm [12] for 10 steps. Each step contains k-space and image correction in sequence, where form the data fidelity and regularizer respectively. XPDNet applies the undersampling mask on the subtraction of the intermediate k-space with original measurements in k-space correction step. Image correction/regularizer is implemented using multi-scale wavelet CNN (MWCNN) [38] followed by a convolutional layer. Inspired by PDNet [1], it uses a modified version of PDHG to utilize a number of optimization parameters instead of just using the previous block’s output. 5 primal and 1 dual variables are used during the unrolling process, and the weights are not shared across the blocks.
RIM implementation based on [67] as described in  [39] unrolls the objective for 16 time steps, where each utilizes a recurrent time step. Each time step takes the previous reconstruction, hidden states and the gradient of negative log-likelihood (as data fidelity term) and outputs the incremental step in image domain to take using a gated recurrent units (GRU) structure [15], where it utilizes depth 1 and 128 hidden channels. Parameters are shared across different recurrent blocks.
E2E-VarNet uses the publicly available implementation [54], and like variational networks, implements an unrolled network to solve the regularized least squares objective using gradient descent. The algorithm is unrolled for 12 steps. Each step combines data fidelity with a regularizer. Data fidelity term applies the undersampling mask after subtraction of intermediate k-space from the measurements, while learned regularizer is implemented via U-Net [74], where it uses 4 number of pull layers and 18 number of output channels after first convolution layer. Weights are not shared across blocks.
Recurrent VarNet uses the publicly available implementation [68] estimates a least squares variational problem by unrolling with gradient descent for 8 steps. Each iteration is a variational block, comprising data fidelity and regularizer terms. Data fidelity term calculates the difference between current level k-space and the measurements on undersampling locations, where regulizer utilizes gated recurrent units (GRU) structure [9]. Each unroll block uses 4 of these GRUs with 128 number of hidden channels for regularizer. Parameters are not shared across different blocks [68].
As described in the main text, all methods were retrained on the respective datasets with supervised learning for maximal performance.
6.2 Comparison Methods and Algorithmic Details
SMUG [37] trains the same PD-DL network we used for MoDL using smoothing via Sec. 2.3. Smoothing is implemented using 10 Monte-Carlo samples [37], with a noise level of $0.01$, where data is normalized in image domain.
Adversarial Training method also uses the same network structure as MoDL. Here, each adversarial sample is generated with 10 iterations of PGD [42] with $\epsilon=0.01$ and $\alpha=\epsilon/5$. Data are normalized to $[0,1]$ in image domain.
Blind Mitigation Schedules. For blind mitigation, our linear scheduler for $\epsilon$ starts from 0.04 and decreases by 0.01 each step until the cyclic loss stabilizes. Then, step size $\alpha$ starts from a large value of $\epsilon$ and gradually decreases, ending at $\epsilon/3.5$ until the cyclic loss shows no further improvement. As mentioned in the main text, since the $\ell_{\infty}$ ball contains the $\ell_{2}$ ball of the same radius, and noting the unitary nature of the Fourier transform in regards to $\ell_{2}$ attack strengths in k-space versus image domain, we always use the $\ell_{\infty}$ ball for blind mitigation.
7 Attack Detection using Simulated k-space
Algorithm 2 Attack Detection
1:${\bf z}_{\Omega},{\bf E}_{\Omega},{\bf E}_{\Delta},f(\cdot,\cdot;{\bm{\theta}}),\tau$triangleright Input parameters
2:True or False, presence of attack $\triangleright$ Output
3:${\bf\tilde{x}}_{\Omega}\leftarrow f({\bf z}_{\Omega},{\bf E}_{\Omega};{\bm{\theta}})$
4:${\bf{y}}_{\Delta_{i}}\leftarrow\mathbf{E}_{\Delta}{\bf\tilde{x}}_{\Omega}+\tilde{\bf n}$
5:${\bf\tilde{x}}_{\Delta}\leftarrow f({\bf E}_{\Delta}^{H}{\bf{y}}_{\Delta},{\bf E}_{\Delta};{\bm{\theta}})$
6:$\zeta_{1}=\frac{||{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf\tilde{x}}_{\Omega}||_{2}}{||{\bf y}_{\Omega}||_{2}}$
7:$\zeta_{2}=\frac{||{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf\tilde{x}}_{\Delta}||_{2}}{||{\bf y}_{\Omega}||_{2}}$
8:If $\zeta_{2}-\zeta_{1}\geq\tau$ True, else False
The description of the attack propagation suggests a methodology for detecting these attacks. Noting that the process is best understood in terms of consistency with acquired data in k-space, we perform detection in k-space instead of attempting to understand the differences between subsequent reconstruction in image domain, which is not clearly characterized. In particular, we define two stages of k-space errors in terms of ${\bf y}_{\Omega}$ for ${\bf\tilde{x}}_{\Omega}$ and ${\bf\tilde{x}}_{\Delta_{i}}$, which were defined in Eq. 8-(9) as follows:
$\zeta_{1}$ $=\frac{||{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf\tilde{x}}_{\Omega}||_{2}}{||{\bf y}_{\Omega}||_{2}},$ (13)
$\zeta_{2}$ $=\frac{||{\bf y}_{\Omega}-{\bf E}_{\Omega}{\bf\tilde{x}}_{\Delta_{i}}||_{2}}{||{\bf y}_{\Omega}||_{2}}.$ (14)
From the previous description $\zeta_{1}$ is expected to be small with or without attack. However, $\zeta_{2}$ is expected to be much larger under the attack, while it should be almost at the same level as $\zeta_{1}$ without an attack. Thus, we check the difference between these two normalized errors, $\zeta_{2}-\zeta_{1}$, and detect an attack if it is greater than a dataset-dependent threshold. The process is depicted in Fig.6, and summarized in Algorithm 2.
Figure 6: The propagation of the attack depicted in Fig.1a suggests a way to track the normalized $\ell_{2}$ error on sampled k-space locations after reconstructions, and a large change in this error is indicative of an attack.
Figure 7 shows how $\zeta_{2}-\zeta_{1}$ changes for knee and brain datasets for both PGD and FGSM attacks on normalized zerofilled images for $\epsilon\in\{0.01,0.02\}$. It is clear that cases with an attack vs. non-perturbed inputs are separated by a dataset-dependent threshold. Note that given the sensitivity of PD-DL networks to SNR and acceleration rate changes, this dataset dependence is not surprising [34], and can be evaluated offline for a given trained model.
Figure 7: Attack detection for different datasets. $\zeta_{2}-\zeta_{1}$ for different attack types are clearly separated from the no attack case. For stronger attack, $\epsilon=0.02$, $\zeta_{2}-\zeta_{1}$ is more easily distinguishable. The violin plots show the median and [25,75] percentile in darker colors for easier visualization.
8 Quantitative Results and Representative Examples
Due to space constraints, the figures and results in the main text focused on $\ell_{\infty}$ attacks generated with unsupervised PGD [42], as mentioned in Sec. 4.2. This supplementary material section provides the corresponding results on related attack types mentioned in Sec. 4.3.
$\epsilon$ Metric SMUG Adversarial Training (AT) Proposed Method + MoDL / SMUG / AT
$0.01$ PSNR $28.22$ $33.99$ $35.14/34.85/\bf{36.57}$
SSIM $0.79$ $0.92$ $0.92/0.92/\bf{0.94}$
$0.02$ PSNR $21.86$ $30.91$ $33.25/32.97/\bf{33.42}$
SSIM $0.61$ $0.88$ $0.91/0.91/{\bf 0.93}$
Table 2: Different attack strengths: Quantitative metrics on all test slices of Cor-PD
8.1 Higher Attack Strengths
Tab. 2 summarizes the quantitative population metrics for different attack strengths, $\epsilon$, complementing the representative examples shown in Fig. 3 of Section Sec. 3.2. These quantitative results align with the visual observations.
8.2 Quantitative Metrics for Different Networks
Tab. 3 shows that the quantitative metrics for the proposed attack mitigation strategy improve substantially compared to the attack for all unrolled networks, aligning with the observations in Fig. 4.
Network Metric With Attack After Proposed Mitigation
XPDNet PSNR $25.49$ $29.43$
SSIM $0.67$ 0.800.80{0.80}0.80
RIM PSNR $19.63$ $34.81$
SSIM $0.39$ $0.90$
E2E-VarNet PSNR $24.24$ 29.5229.52{29.52}29.52
SSIM $0.59$ $0.84$
Recurrent VarNet PSNR $22.27$ 29.2429.24{29.24}29.24
SSIM $0.52$ $0.84$
Table 3: Quantitative metrics for different unrolled networks
8.3 Different Adversarial Training Methods
This subsection provides an alternative implementation of the adversarial training based on Eq. 6 with $\lambda=1$ to balance the perturbed and clean input, instead of Eq. 5 that was provided in the main text as a comparison. Results in Tab. 4 show that the version in the main text outperforms the alternative version provided here.
Method Metric With Attack
Adversarial Training (AT) with Eq. (5) PSNR 33.9933.99{33.99}33.99
SSIM 0.920.92{0.92}0.92
AT with Eq. (6) PSNR $33.61$
SSIM $0.91$
AT with Eq. (5) + Proposed Method PSNR $36.17$
SSIM 0.940.94{0.94}0.94
AT with Eq. (6) + Proposed Method PSNR 36.9136.91{36.91}36.91
SSIM 0.940.94{0.94}0.94
Table 4: Comparison of adversarial training approaches.
8.4 Mitigation Performance on Non-Perturbed Data
As discussed in Sec. 3.1, Algorithm 1 does not compromise the reconstruction quality if the input is unperturbed. This is because, with an unperturbed input image in Sec. 3.2, the intermediate reconstruction remains consistent with the measurements. As a result, the objective value remains close to zero and stays near that level until the end, indicating the mitigation starts from an almost optimal point of the objective function. Hence, the mitigation does not degrade the quality of the clean inputs, and does not incur large computational costs, as it effectively converges in a single iteration. Visual examples of this process are depicted in Fig. 8.
Figure 8: Performance of mitigation algorithm on non-perturbed data. The mitigation effectively converges in one iteration. As it shown, the algorithm maintains the quality of the clear input.
Figure 9: Performance of different methods under FGSM attack.
8.5 Supervised Attacks
While Sec. 4.2 and 4.3 focused on unsupervised attacks due to practicality, here we provide additional experiments with supervised attacks, even though they are not realistic for MRI reconstruction systems. Tab. 5 shows that the proposed method is equally efficient in mitigating supervised attacks.
Attack Method Metric Proposed Method
Unsupervised Attack PSNR $32.44$
SSIM ${\bf 0.91}$
Supervised Attack PSNR ${\bf 32.55}$
SSIM ${\bf 0.91}$
Table 5: Mitigation with Supervised vs. Unsupervised Attacks
8.6 FGSM Attack
In Sec. 4.2, we used the PGD method for attack generation due to the more severe nature of the attacks. Here, we provide additional experiments with FGSM attacks [19]. Tab. 6 show results using SMUG, adversarial training and our method with FGSM attacks with $\epsilon=0.01$. Corresponding visual examples are depicted in Fig. 9 , showing that all methods perform better under FGSM compared to PGD attacks.
Metric SMUG Adversarial Training (AT) Proposed Method + MoDL / SMUG / AT
PSNR ${\bf 36.24}$ $35.61$ ${\bf 36.24}$ / $35.13/36.06$
SSIM ${\bf 0.93}$ ${\bf 0.93}$ ${\bf 0.93}$ / $0.92/\bf{0.93}$
Table 6: FGSM attack: Quantitative metrics on all test slices of Ax-FLAIR
8.7 $\ell_{2}$ Attacks in k-space
$\ell_{2}$ attacks have been used in k-space due to the large variation in intensities in the Fourier domain [50]. To complement the $\ell_{\infty}$ attacks in image domain that was provided in the main text, here we provide results for $\ell_{2}$ attacks in k-space, generated using PGD [42] for 5 iterations, with $\epsilon=0.05\cdot||{\bf y}_{\Omega}||_{2}$ and $\alpha=\frac{\epsilon}{5}$. Fig. 10 depicts representative reconstructions with $\ell_{2}$ attacks in k-space using baseline MoDL, adversarial training and our proposed mitigation. Table 7 shows comparison of adversarial training and the proposed method on Cor-PD datasets, highlighting the efficacy of our method in this setup as well.
We also emphasize that the $\ell_{\infty}$ image domain attacks are easily converted to attacks in k-space, which are non-zero only on indices specified by $\Omega$, as described in Sec. 2.2.
Method Metric $\ell_{2}$ Attack
Adversarial Training PSNR $33.37$
SSIM $0.88$
Proposed Method + MoDL PSNR ${\bf 34.21}$
SSIM ${\bf 0.89}$
Table 7: Mitigation results for $\ell_{2}$ attacks in k-space
8.8 Further Blind Mitigation
Here, we provide results for using blind mitigation with $\ell_{2}$ attacks in k-space. Fig. 11 depicts example reconstructions with $\ell_{2}$ attacks in k-space using baseline MoDL and our blind mitigation approach. Tab. 8 compares our blind mitigation approach to our mitigation strategy with known attack type and level, showing that blind mitigation performs on-par with the latter for both $\ell_{2}$ attacks in k-space and $\ell_{\infty}$ attacks in image domain.
Attack Method Metric Proposed Method ($\ell_{\infty}$ attack) Proposed Method ($\ell_{2}$ attack)
Knowing the Attack PSNR ${\bf 33.23}$ ${\bf 34.21}$
SSIM ${\bf 0.92}$ ${\bf 0.89}$
Blind Mitigation PSNR 32.9432.94{32.94}32.94 $33.73$
SSIM ${\bf 0.92}$ $0.88$
Table 8: Blind mitigation for $\ell_{2}$ (k-space, $\epsilon=0.05\cdot||{\bf y}_{\Omega}||_{2}$) and $\ell_{\infty}$ (image domain, $\epsilon=0.01$) attacks.
Figure 10: Representative reconstructions under $\ell_{2}$ attack on measurements with $\epsilon=0.05\cdot||{\bf y}_{\Omega}||_{2}$ using MoDL, adversarial training, and our proposed method.
8.9 Non-Uniform Undersampling Patterns
While the main text focused on uniform undersampling, which is considered to be a harder problem [24, 61], here we describe results with random undersampling, generated with a variable density Gaussian pattern [2]. All networks were retrained for such undersampling patterns. The attack generation and our mitigation algorithms were applied without any changes, as described in the main text. Fig. 12 shows representative examples for different methods, highlighting that our method readily extends to non-uniform undersampling patterns. Tab. 9 summarizes the quantitative metrics for this case, showing that the proposed mitigation improves upon MoDL or adversarial training alone.
Metric MoDL Adversarial Training (AT) Proposed Method + MoDL / AT
PSNR $22.30$ $32.22$ $31.82/{\bf 34.12}$
SSIM $0.62$ $0.89$ $0.87$ /${\bf 0.92}$
Table 9: Attacks on non-uniform undersampling
Figure 11: Representative reconstructions under $\ell_{2}$ attack using MoDL and our proposed blind mitigation.
Figure 12: Representative reconstructions for non-uniform undersampling reconstructions using MoDL, adversarial training, and our proposed method under adversarial attacks.
9 Further Details on Adaptive Attacks
This section contains more information about adaptive attack generation and visual examples.
9.1 Hyperparameter Tuning for Adaptive Attacks
The parameter $\lambda$ in Sec. 3.3 balances the two terms involved in the adaptive attack generation. A higher $\lambda$ produces a perturbation with more focus on bypassing the defense strategy, while potentially not generating a strong enough attack for the baseline. Conversely, a small $\lambda$ may not lead to sufficient adaptivity in the attack generation. To this end, we computed the population-average PSNRs of the reconstruction after the iterative mitigation algorithm on a subset of Cor-PD for various $\lambda$ values for $T\in\{10,25\}$, as shown in Tab. 10. These results show that $\lambda=5$ leads to the most destructive attack against our mitigation algorithm, and was subsequently used for adaptive attack generation in Sec. 4.3.
Unrolls $\lambda=1$ $\lambda=2$ $\lambda=3$ $\lambda=5$ $\lambda=10$
$T=10$ $34.51$ $34.48$ $34.41$ ${\bf 34.34}$ $34.66$
$T=25$ $34.41$ $34.36$ $34.40$ ${\bf 34.16}$ $34.47$
Table 10: Fine tuning the $\lambda$ parameter in Sec. 3.3 across $T\in\{10,25\}$
9.2 Verification of Gradient Obfuscation Avoidance
While our adaptive attack implements the exact gradient to avoid gradient obfuscation (including shattered, stochastic, and vanishing gradients [7]), there are some methods to verify that gradients are indeed not obfuscated [7]. In particular, we tested two well-established key criteria: 1) One-step attacks should not outperform iterative-based ones, and 2) Increasing the perturbation bound (i.e. $\epsilon$) should lead to a greater disruption. Tab. 11 summarizes these two criteria, showing PSNRs of the iterative mitigation algorithm output. These demonstrate that a single-step attack cannot surpass the iterative-based ones in terms of attack success, and similarly, increasing the perturbation bound leads to more severe degradation with PGD. These sanity checks align with the fact that we used the exact gradient through the steps described in Sec. 4.2, validating that gradient obfuscation did not happen in our implementation.
$T$ PGD ($\epsilon=0.01$) PGD ($\epsilon=0.02$) FGSM ($\epsilon=0.01$)
10 34.34 33.11 35.17
25 34.16 32.77 35.01
Table 11: Checking Gradient Obfuscation on Cor-PD Dataset over $T\in\{10,25\}$
Figure 13: Representative examples of the mitigation algorithm outputs for adaptive attacks. The number of unrolls $T\in\{10,25,50,100\}$ specified for each adaptive attack on the top. The top row is the baseline reconstruction, where the non-adaptive attack shows more artifacts than adaptive ones, as expected. The second row shows the mitigation outputs using the unrolled version of Algorithm 1, where the number of unrolls are matched between the adaptive attack generation and mitigation. At smaller $T$ values, the unrolled mitigation suffers from performance degradation. Finally, the last row shows the results of the iterative mitigation algorithm on the adaptive attacks. Iterative mitigation, when run until convergence, resolves the attacks, albeit with a slight degradation for high $T$ values. This is consistent with its physics-based design, showing its robustness to adaptive attacks.
9.3 Visualization of Adaptive Attacks and Mitigation
Representative examples showing the performance of the mitigation algorithm for different adaptive attacks generated using Sec. 3.3 with the unrolled version of $g(\cdot;\cdot)$ for $T\in\{10,25,50,100\}$ are provided in Fig. 13. The first row shows the results of the baseline reconstruction under both non-adaptive and adaptive attacks for various $T$. Consistent with Tab. 2, as $T$ increases for the adaptive attack, the baseline deterioration becomes less substantial. The second row shows the performance of the mitigation algorithm when it is unrolled for the same number of $T$ as in the adaptive attack generation. In this case, for lower $T$, the unrolled mitigation has performance degradation, as expected. Finally, the final row shows results of the iterative mitigation algorithm run until convergence. In all cases, the iterative mitigation algorithm successfully recovers a clean image, owing to its physics-based nature, as discussed in Sec. 4.3. However, we note that though adaptive attacks have milder effect on the baseline with increasing $T$, they do deteriorate the iterative mitigation albeit slightly as a function of increasing $T$.
10 Ablation Study
As discussed in Sec. 4.4, we analyzed the number of reconstruction stages for mitigation. By extending the number of reconstruction stages, we can reformulate this by updating the second term in the loss function in Sec. 3.2 to include more reconstruction stages, for instance with 3 cyclic stages instead of 2 given in Sec. 3.2:
$\arg\min_{{\bf r^{\prime}}:||{\bf r^{\prime}}||_{p}\leq\epsilon}\mathbb{E}_{\Gamma}\mathbb{E}_{\Delta}\Bigg{[}\bigg{\|}({\bf E}_{\Omega}^{H})^{\dagger}({\bf z}_{\Omega}+{\bf r^{\prime}})-$
${\bf E}_{\Omega}f\bigg{(}{\bf E}_{\Gamma}^{H}\Big{(}{\bf E}_{\Gamma}f\big{(}{\bf E}_{\Delta}^{H}\big{(}{\bf E}_{\Delta}f({\bf z}_{\Omega}+{\bf r^{\prime}},{\bf E}_{\Omega};{\bm{\theta}})$
$\qquad\qquad\qquad+{\bf\tilde{n}})\big{)},{\bf E}_{\Delta};{\bm{\theta}}\Big{)}+{\bf\tilde{n}},{\bf E}_{\Gamma}\Big{)};{\bm{\theta}}\bigg{)}\bigg{\|}_{2}\Bigg{]}.$ (15)
Empirically, in our implementation, we carry out the expectation over all possible permutations without repeating any patterns. As a result, the error propagated to the last stage becomes larger, as we rely more on synthesized data. In turn, this makes the optimization process harder, deteriorating the results, as shown in Fig. 14. Consequently, in addition to these performance issues, the computation costs of adding more cyclic reconstruction is often impractical, leading to the conclusion that 2-cyclic stages as in Sec. 3.2 are sufficient.
Figure 14: Ablation study on the number of stages for cyclic measurement consistency shows that 2 levels of reconstruction (left) is better than more levels (middle, right), as the latter has stronger reliance on synthesized k-space data.